Jane wishes to have $20,000 available at the end of 10 years, so she deposits money into an account that pays 1.14% compounded monthly. How much does she need to deposit in order to meet this goal?
Answer #1
Jane wishes to have $ 20,000 at end of 10 Years. So,
Future Value = 20000
Interest rate compounded monthly (r )= 1.14% or 0.0114
This rate is annual rate compounded monthly.
So, Effective interest rate per year = (1+ r/m)m - 1
Effective Annual interest rate = ( (1 + 0.0114/12) 12 } - 1
Effective Annual interest rate = 0.01146
Effective Monthly Interest rate = 0.01146/12 = 0.000955
Time period = 10 years
Compounding is monthly. So total monthly periods = 10 *12 = 120
Present Value = Future Value / (1+r)n
.           = 20000 / (1+0.000955)120
$17,835.43
So, she need to deposit $ 17,835.43 to have $ 20,000 at end of 10 Years.